Present value of annual inflows (3,024,000) discounted at 6% for 25 years = $24,682,289.17

how do you find this PV by hand, not calculator?
Answer #1

PV annuity formula =(1-(1+r)^(-n))/r

r=interest Rate

n- number of period

PV =3024000*(1-(1+6%)^(-25))/.06=$38,656,869.02